OpenStudy (anonymous):

3x − 2 = 5x − 6 solve please

OpenStudy (unklerhaukus):

add six to both sides of the equation

OpenStudy (unklerhaukus):

and take away 3x from both sides of the equantion

OpenStudy (anonymous):

Which one is correct

OpenStudy (unklerhaukus):

3x - 2 = 5x - 6 well , you could add two, to both sides of the equation 3x - 2 + 2 = 5x - 6 + 2

OpenStudy (anonymous):

So the left one is correct? where did Aaron mess up

OpenStudy (unklerhaukus):

the mistake is made when one of them performs an operation of one side of the equation, but not the other side as well

OpenStudy (unklerhaukus):

(the operations are marked in red)

OpenStudy (anonymous):

ahhh so when aaron subtracted -6

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

yeah!, what should he have gotten for that line?

OpenStudy (unklerhaukus):

-2 = 2x - 6 Aaron should have added six, and to both sides -2 + 6 = 2x - 6 + 6 4 = 2x